can someone help me plss before monday?
solve the equation.
19. 2x/x-6 - 1/x+6 = 27/x^2-36

Answer by jorel1380(2518)

2x/x-6 - 1/x+6 = 27/x^2-36
2x(x+6)/(x2-36)-1(x-6)/(x2-36)=27/x2-36
2x2+12x-x+6=27
2x2+11x-21=0
(2x-3)(x+7)=0
x=3/2 or -7..